OpenStudy (anonymous):

Convert 132 F TO Celsius. 55 C 70 C 85 C 100 C

OpenStudy (nurali):

55.5556

OpenStudy (anonymous):

(132-32)*5/9

OpenStudy (juana02):

Exactly 70 °F is equal to about 21.1 °C The conversion formula is : °C = (°F - 32) x 5/9 = (70 - 32) x 5/9 = 21.1111 degrees Celsius

OpenStudy (anonymous):

55 your answer is a

OpenStudy (juana02):

celsius = (fahrenheit - 32) x 5/9 = (50-32) x 5/9 = 18 x 5/9 = 10 degrees Celsius

OpenStudy (juana02):

Exactly 85 °F is equal to 29.44 °C The conversion formula is Celsius temperature = 5/9 x (Fahrenheit temperature - 32) Start by taking the number in Fahrenheit and subtracting 32. Then multiply the number by 5, and then divide by 9. 85 - 32 = 53 53 x 5 = 265 265 / 9 = about 29.44 It is about 29.44

OpenStudy (juana02):

100 °F = 37.78 °C The conversion formula is Celsius temperature = 5/9 x (Fahrenheit temperature - 32) 5/9 x (100-32) = 5/9 (68) = 37.78